In a city, it is estimated that the maximum temperature in June is

normally distributed with a mean of 23o


and a standard deviation of 5

o

.

Calculate the number of days in this month in which it is expected to

reach a maximum of between 21° and 27°.


1
Expert's answer
2021-05-07T14:21:25-0400

"P(21<X<27)=P(\frac{21-23}{5}<Z<\frac{27-23}{5})=P(-0.4<Z<0.8)="

"=P(Z<0.8)-P(Z<-0.4)=0.4435."

Number of days: "n=30*0.4435=13."